January field season update

Hi Everyone,

This update is a group effort, mostly put together by Diana, and sent by Kathy.  Today is supposed to be Diana's last day as the MCM-LTER on ice POC. 

A flurry of work is going on as we approach the end of the season. The ship offload folks are in town and the town population is rising. Weather in MCM was a little foggy with no fly days on Monday, Wed and partial on Tuesday and as typical, weather was terrific in the valleys. By the end of the week, there was a lot of flying and lots of work accomplished thanks to Lindsay and Kate who did an awesome job of scheduling and the pilots who were supremely generous of their time.   


504-Glacier Team:

Anna headed home after a successful season. Adam and Forest were spotted hoofing it all over the glaciers checking met stations and ablation stakes.

505-Limno team: Amy has been approved to stay an extra 5 days, and will be departing on 30 Jan. She has finished the Limno DIC samples, will be working on radioisotope checkout and other miscellaneous tasks this week, and will also be collecting a Lake Chad water sample on Wednesday.

506- Stream team:  The stream team continues their stream monitoring and will be returning to McMurdo on 30 Jan. Algae Ops continues with Aneliya, who will be returning to McMurdo this week to run chlorophyll samples.

507- The Wormherder Soils team has been on the move: Tandra and Jeb left the ice Thursday; Ruth and Jess completed the CO2 at SSLB and F6 and got to Miers before coming back to town. Byron and Dan took a day sampling the BSNEs. Ashley, Summer and Byron got the batteries changed out for Miers and SSLH soil met Stations and sampled Beacon Valley for protists; Diana, Ashley, and Summer were at Parera and Marr Pond where they were met by Tom Wilch and some DVs.  Afterwards Tom stopped by and said the DVs (4-star generals et al) made multiple references to how impressed they were with our students. Ruth has led the whole team in getting the chlorophyll a’s completed, and Walter has been central to operations. Close out will be Thursday.

509 – Geochem team:  Kathy, Elsa and Sydney are making progress on analytical work and completed packing of samples for the vessel deadline. They made a day trip to Lower Miers on Friday. 

 Rae and Renee had a visit from the DVs and word traveled fast in the field about ‘DV cookies”, and for those who got to taste some of those cookies, awesome!


Bad news:

Byron claimed he could smell the Lake Hoare cookies all the way from the worm farm on the other side of the lake.  That’s not so bad, but he said it would have been even better to eat a few of them.


The BSNEs (aeolian collectors) had a rough year (see photo).  When Byron went to sample them, he found the ones at F6, Taylor Glacier, Howard Glacier and Commonwealth Glacier had blown over.  The ones on the glaciers had melted in, and then were snowed over.  But they made for some funky shaped cryoconites (see photo).  Byron was able to put the ones at Commonwealth Gl. and F6 back into service, but retro’d the ones on the Taylor and Howard Gl.  They’ve been repaired and can probably be redeployed next season.

 All of us here send our regards and best wishes.

:993/fetch%3EUID%3E.INBOX%3E115833?part=1.2&type=image/jpeg&filename=collector.jpg" />

Article Section: